### v1

# Studying Ionospheric Phase Structure Functions Using Wide-Band uGMRT (Band-4) Interferometric Data

Abstract
Interferometric observations of the low-frequency radio sky ($<$ 1 GHz) are largely limited by systematic effects introduced by the ionosphere. In this study, we use a ten-hour nighttime observation of the bright radio source 3C48, carried out with the upgraded Giant Metrewave Radio Telescope (uGMRT), to examine how ionospheric conditions affect radio signals. We focus on measuring phase fluctuations caused by the ionosphere and analyse how these variations change with antenna separation using the spatial phase structure function. Our results, based on data from three sub-bands between 575 and 725 MHz, show a clear power-law trend consistent with turbulent behaviour. We introduce the diffractive scale as a scalar measure of ionospheric conditions, which could guide future calibration strategies. Furthermore, we find that the turbulence is anisotropic, with properties that vary by direction. The smallest diffractive scales do not always align with Earth’s magnetic field, suggesting the influence of medium-scale Travelling Ionospheric Disturbances (MSTIDs). The contribution of these wave-like MSTIDs to the structure function follows a power-law with a slope close to 2.0, indicating their dominant role in shaping the observed anisotropic phase fluctuations.
keywords:
methods - statistical, data analysis, techniques - interferometric, atmospheric effects – instrumentation
1 Introduction
At low radio frequencies, the ionosphere is the main contributor to phase disruptions, affecting signals from astronomical sources (Taylor et al., 1999). A monochromatic interferometer detects the phase shift caused by differences in signal delay as the wavefront passes through the ionosphere at two separate locations. These phase shifts are due to variations in the mean refractive index along the signal paths reaching each antenna.
As long as propagation remains in the refractive regime, the interferometer’s phase differences can be interpreted as differences in excess path length the additional distance light appears to travel due to the ionosphere between the two ray paths. This assumes that signals from the source travel nearly parallel through the ionosphere to each antenna. In the diffractive regime, however, wavefronts lose their parallel alignment, and intensity fluctuations can arise not only between antennas but even across a single antenna’s surface. In such cases, calibration methods based on refractive assumptions break down (Vedantham and Koopmans, 2015; de Gasperin et al., 2019).
The excess path length results from the deviation of the ionospheric refractive index from unity, integrated along the geometric signal path. This effect depends on the column density of free electrons, known as the Total Electron Content (TEC), along the line of sight.
A useful statistical tool for characterizing ionospheric phase irregularities is the spatial phase structure function, which describes how phase fluctuations vary with antenna separation across an array. It quantifies the average squared difference in interferometric phase between pairs of antennas as a function of their relative spacing. This function captures how phase variability changes with scale and provides insight into the spatial behaviour of ionospheric disturbances.
In turbulent media like the ionosphere, the structure function often follows a power-law relationship with antenna separation. The exponent in this power law, typically denoted as $\beta$, reveals the nature of the underlying turbulence. A value of $\beta=5/3$ is expected for fully developed isotropic turbulence, following Kolmogorov theory. However, deviations from this value can indicate the presence of other effects, such as organized wave structures or anisotropic conditions in the ionosphere (Spencer, 1955; Mevius et al., 2016; Jordan et al., 2017; Loi et al., 2015).
Although ionosondes and GPS-based methods provide valuable measurements of large-scale ionospheric properties (Arora et al., 2016; de Gasperin et al., 2018; Fallows et al., 2020), radio interferometers offer unprecedented resolution for probing ionospheric structures on finer spatial and temporal scales. Arrays such as Low-Frequency Array (LOFAR), Murchison Widefield Array (MWA), Very Large Array (VLA), and more recently, Square Kilometre Array (SKA) pathfinders, have been used to examine ionospheric behaviour with high precision. These studies have revealed rich ionospheric dynamics, including anisotropy aligned with the Earth’s magnetic field, spectral signatures of TIDs, and turbulent cascades across a wide range of spatial scales (Rufenach, 1972; Vanvelthoven, 1990; Helmboldt et al., 2012b, a; Loi et al., 2015).
The upgraded Giant Metrewave Radio Telescope (uGMRT), situated near Pune in India is a powerful but still underused tool for studying the ionosphere (Swarup et al., 1991; Gupta et al., 2017). It has thirty large 45-meter dishes and can observe a broad spectrum of radio frequencies (from 120 to 1460 MHz) using five different frequency bands. Thanks to its improved receivers and advanced processing systems, the uGMRT can measure ionospheric phase changes with high sensitivity and detail across both time and frequency. Its location in a mid-latitude region is also important, as it allows scientists to study ionospheric conditions that are different from those found near the poles or the equator (Mangla and Datta, 2022, 2023).
In this study, we take a close look at the ionospheric phase structure function derived from a single ten-hour Band-4 observation conducted with the upgraded Giant Metrewave Radio Telescope (uGMRT; (Gupta et al., 2017)). The observation targets one of the brightest flux calibrators in the sky, 3C48. 3C48 is a distant quasar outside our galaxy that is often used in radio astronomy to calibrate signal strength and correct for frequency-related effects (Perley and Butler, 2013). It is compact and very bright, with a stable signal across a broad range of radio frequencies, which makes it especially useful for studying ionospheric effects.
The paper is structured as follows: Section 2 outlines the observational data and the processing of phase information. Section 3 presents the concept of ionospheric spatial variability and introduces the two-dimensional structure function employed for its analysis. Here, we present results from modelling both 1D and 2D structure functions across three 25 MHz sub-bands. Additionally, by comparing the orientation of baseline vectors with the local projected magnetic field direction at each time step, we investigate anisotropy in ionospheric phase fluctuations. This analysis helps us understand how ionospheric turbulence aligns with geomagnetic field lines, a phenomenon previously observed in LOFAR (Mevius et al., 2016) and MWA (Loi et al., 2015) studies. Finally, Section 4 represents our discussion and conclusions.
2 Observation and Processing
The Giant Metrewave Radio Telescope (GMRT) is a sensitive radio interferometer optimized for low-frequency observations. It comprises 30 parabolic antennas, each with a diameter of 45 meters, distributed over a 25 km region. Fourteen antennas are located within a central $1.4\times 1.4$ km2 square, and the remaining 16 are spread along three 14 km-long arms forming a ‘Y’-shaped layout. This configuration enables high-resolution imaging and multi-directional probing of ionospheric irregularities. Located between the magnetic equator and the northern crest of the equatorial ionization anomaly (EIA), the GMRT is ideally positioned for ionospheric studies (Mangla and Datta, 2022).
We conducted 10 hours of nighttime observations of the bright quasar 3C48 using the upgraded GMRT (uGMRT) at Band-4 (project code: $47\_003$  [^1] https://naps.ncra.tifr.res.in/goa/data/search) in November 2024. The data span a 200 MHz bandwidth (550–750 MHz) with a central frequency of 650 MHz, using 8192 channels at 24.41 kHz resolution and an integration time of 2.68 seconds. For ease of computation and to avoid frequency ranges significantly affected by radio frequency interference (RFI), we selected a subset of three relatively clean sub-bands: 575–600 MHz, 600–625 MHz, and 700–725 MHz. We found overall good agreement among the analysed bands. More details about the observations are summarized in Table 1.
[…]
Table 1: Observation summary
Observation ID 47_003
Observation Date 23 November 2024
Observation Window (IST) 5:21 PM to 3:27 AM
Processed Frequency Bands 575–600, 600–625, and 700–725 MHz
Channel Width 24.41 kHz
Integration Time 2.68 seconds
Measured Correlations RR and LL
Functional Antennas 30
Target Source 3C48
Right Ascension (J2000) 01h37m41.301596s
Declination (J2000) +330935.25459′′
Total Exposure Time 10 h
2.1 Calibration and Phase Processing
After RFI removal, direction-independent calibration was performed using standard CASA tasks with 3C48 as the flux, bandpass, and phase calibrator. Flux density was set using the Scaife and Heald (2012) model via the CASA task SETJY. Calibration began with delay corrections using GAINCAL (gaintype=K) relative to a reference antenna (C06), followed by bandpass calibration using BANDPASS with a solution interval of 10 seconds and 8 channels ($\sim$194 kHz). Although both amplitude and phase are calibrated, we focus here on the phase solutions for ionospheric studies. A total of $\sim 32\%$ of the data was flagged post-calibration and RFI mitigation.
Phase differences between antennas were modelled as contributions from ionospheric, instrumental, source structure, and phase ambiguity effects (Helmboldt et al., 2012b; Mangla and Datta, 2022). Source-related phase contributions were removed using model visibilities, while $2\pi$ ambiguities were addressed through phase unwrapping. Outliers and phase jumps were detected and flagged using a Local Outlier Factor (LOF) algorithm combined with a cosine filter (Banerjee et al., 2025). Time steps with abrupt phase changes (exceeding $5\sigma$) were flagged using a cosine filter (Helmboldt et al., 2012b) and interpolated to maintain temporal continuity. Further details regarding the phase correction process can be found in Banerjee et al. 2025.
To isolate ionospheric fluctuations from slowly varying instrumental phases, we applied a one-hour boxcar smoothing filter to the unwrapped phases, treating the residual as ionospheric variability. Although absolute instrumental phases could not be fully separated, this approach enabled analysis of short-timescale TEC variations ($\leq$ 1 hour). Residual phase variations post-continuum subtraction were typically within ±$\pm 0.5$ radians for central square antennas and up to ±$\pm 5.0$ radians for arm antennas (Banerjee et al., 2025). A notable increase in residual scatter between 2-4 hours suggests enhanced ionospheric activity or unflagged RFI, consistent with scintillation reports in the uGMRT observation log. These fluctuations were significantly stronger than those in adjacent time steps, leading us to consider the data unreliable and flag this time interval (Banerjee et al., 2025).
3 Ionospheric Phase Structure Function
The phase structure function is a key tool in characterizing ionospheric turbulence. It encapsulates the scale-dependent nature of phase fluctuations due to turbulent plasma density variations. For Kolmogorov-type turbulence, both the phase and refractive index structure functions exhibit power-law behaviour, and are tightly linked to their respective power spectra. These formulations are essential in modelling ionospheric effects on radio wave propagation, particularly for applications in radio astronomy, satellite communication, and GNSS positioning (Van der Tol, 2009).
The spatial characteristics of a turbulent medium like the ionosphere are usually described using its power spectrum, or alternatively its phase correlation function, which corresponds to the Fourier transform of the power spectrum. However, in practice, it is often more practical to describe how the ionospheric phase changes across space using the phase structure function (Mevius et al., 2016),
$\Xi(r)=\left\langle\left(\phi(r^{\prime})-\phi(r^{\prime}+r)\right)^{2}\right\rangle,$ (1)
where $\phi(r)$ represents the ionospheric phase at position $r$, and the angle brackets denote an ensemble average. This function measures the expected squared difference in phase between two antenna locations spaced by a distance $r$.
In the case of Kolmogorov turbulence, which describes a wide range of turbulent media including the ionosphere, the phase structure function follows a power-law behaviour within the inertial range:
$\Xi(r)=\left(\frac{r}{r_{\rm diff}}\right)^{\beta},$ (2)
where $r_{\rm diff}$ denotes the diffractive scale, the spatial separation at which the phase variance reaches $1~{}\mathrm{rad}^{2}$, and $\beta=5/3$ corresponds to pure Kolmogorov turbulence (Narayan, 1992). This scale encapsulates the strength of the turbulence: smaller values of $r_{\rm diff}$ indicate stronger phase fluctuations over smaller spatial separations.
Figure 1: Differential TEC ($\delta{\rm TEC}$) for the RR polarization of uGMRT baselines relative to the reference antenna ‘C06’ as a function of time. The colour bar indicates baseline length. (Top row) Baselines within the central square. (Bottom row) Extended antenna arm baselines.
To estimate the structure function using uGMRT B4 data, we compute the variance of the differences in Total Electron Content (TEC) for each baseline, as shown in Figure 1. We notice due to the large-scale structure of the ionosphere, such as Travelling Ionospheric Disturbances (TIDs), the average differential TEC ($\delta{\rm TEC}$) between two antenna is not zero. We convert the slant TEC (measured along the line of sight) into vertical TEC by correcting for the angle of observation (Helmboldt et al., 2012b; Banerjee et al., 2025). We assume the ionosphere is a single thin layer with the peak height is predicted by the IRI model (https://www.ionolab.org/iriplasonline/) during the time of our observation. We then convert the vertical differential TEC into phase delay at 600 MHz, using Equation (3), to compare with other experiments.
$\Delta\phi_{\text{ion}}\approx 8.45\left(\frac{1\,\text{GHz}}{\nu}\right)\left(\frac{\delta\text{TEC}}{1\,\text{TECU}}\right)\quad\text{radians}$ (3)
During our observation, we track a bright radio source 3C48, which effectively means observing the ionosphere across a projected path of about 300 km (assuming the thin layer of ionosphere is at an average height of $\sim$ 300 km). At the same time, the ionosphere itself is moving — for example, Travelling Ionospheric Disturbances (TIDs) can move at several hundred kilometres per hour. This movement increases the area we sample, although in some cases the movement direction of the ionosphere and the source can align.
Because of the long observation time and movement, we collect enough independent measurements of the ionosphere to assume that statistical averages (ergodic theorem) apply. Figure 2 also shows an example of the spatial un-binned phase structure function for a typical Band-4 uGMRT night-time observation.
Figure 2: Phase structure function at approximately 587.5 MHz. The blue points represent the measured phase variance as a function of baseline length, while the red solid line shows the fitted one-dimensional power-law model.
The structure function displays a clear power-law trend across a broad range of baseline lengths, from roughly 100 meters up to 25 kilometres. There is no sign of it flattening out at the longest baselines (around 20 km), which would typically suggest the outer scale of turbulence - where the structure function is anticipated to stop increasing and become constant. To determine the outer scale of turbulence more definitively, we would need to include much longer baselines (Mevius et al., 2016), which we currently don’t have access to. There is ongoing effort to extend uGMRT baselines upto 100 km (Patra et al., 2019), this can enhance the angular resolution of the GMRT by a factor of 5 and also help us to determine the outer scale of turbulence.
Figure 2 shows the measured phase structure function along with the best-fit 1D model. The results follow a power-law with an index of $\beta=1.71\pm 0.07$, which is slightly steeper than the expected Kolmogorov value of $1.66$. This suggests that there may be additional structures in the ionosphere affecting the measurements (Vanvelthoven, 1990; Loi et al., 2015). The fitted diffractive scale is $r_{\rm diff}=6.68$ km, which is the distance where the phase variance reaches 1 rad2.
A small diffractive scale means that the phase of the signal changes a lot across the field of view. When the diffractive scale is small, the signal also stays coherent for a shorter time, and this depends on the length of the baseline (Vedantham and Koopmans, 2015). To correct for these fast-changing phase variations, calibration must be done in many directions and at short time intervals (de Gasperin et al., 2019). However, the number of directions and how often we can update the calibration depends on how bright the sources are and how much independent data is available (Intema, H. T. et al., 2009; Bregman, 2012). Overall, $r_{\rm diff}$ is a useful indicator of how stable the ionosphere is for radio imaging during the night.
At short baseline lengths, the structure function can also flatten. This effect has been observed in a recent LOFAR study of 29 nights of data from the 3C196 field, and is primarily caused by a noise floor (Mevius et al., 2016). Small calibration inaccuracies arising from an incomplete sky model or beam model errors can introduce second-order effects that are not accounted for during the clock/TEC separation process, these factors cause small but consistent variations in the estimated TEC values.
In our uGMRT observations, we do not observe such flattening. However, if we assume that any potential flattening at short baselines is entirely due to uncorrelated systematics. This can be modelled by incorporating a constant noise term into the phase structure function, which can then be used to estimate the uncertainty in the TEC values. In this work, however, we restrict ourselves to a pure power-law structure function model, without any additive noise term.
3.1 2D Structure Function
In Figure 2, we notice a band-like pattern in the phase structure function. This pattern seems to depend on the direction of the baseline. This kind of band-like structure was also earlier reported in LOFAR observations (Mevius et al., 2016). We find that the scale of the structure function (i.e., the $r_{\text{diff}}$ value) varies with direction, even though the slope remains roughly the same. This indicates that ionospheric irregularities are anisotropic, meaning their properties differ depending on direction.
The observed banded pattern may also be attributed to the presence of large, wave-like structures within the ionosphere. These waves appear to propagate in the direction where the smallest diffractive scales are found. The anisotropy of the ionosphere has been reported in multiple earlier studies (Wheelon, 2001; Spencer, 1955; Singleton, 1970; Mevius et al., 2016).
To account for this anisotropy, the structure function can be generalized to a two-dimensional form that allows for direction-dependent variations (Figure 3).
Figure 3: Two-dimensional phase structure function at 587.5 MHz. The green dotted and red solid lines represent the major and minor axis projections, respectively, based on the 2D structure function model described in Equation 4.
The anisotropic structure function is defined as,
$\Xi(\mathbf{r})=\left(\mathbf{r}^{\top}R^{\top}CR\mathbf{r}\right)^{\beta/2},$ (4)
where $R$ denotes the $2\times 2$ rotation matrix that aligns the coordinate system with the principal axes of anisotropy:
$R=\begin{bmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{bmatrix},\quad C=\mathrm{diag}\left(\frac{1}{r_{\mathrm{diff,maj}}^{2}},\frac{1}{r_{\mathrm{diff,min}}^{2}}\right).$
Here, $\mathbf{r}=(r_{x},r_{y})$ be the baseline vector, where $r_{x}$ and $r_{y}$ are its components along the east-west (EW) and north-south (NS) directions, respectively. The diffractive scales $r_{\mathrm{diff,maj}}$ and $r_{\mathrm{diff,min}}$ are defined with respect to an orthogonal coordinate system aligned with these directions. $\alpha$ is the angle of the major axis measured from geographic east, and $\beta$ is the power-law exponent as in the isotropic case.
We calculated the phase structure function and fitted it using the 2D model (equation (4)) to estimate the parameters $r_{\mathrm{diff}}^{\mathrm{maj}},r_{\mathrm{diff}}^{\mathrm{min}},\alpha,{\rm and}\beta$, which describe the ionospheric conditions. For consistency and improved stability of the fit, the value of $\beta$ was fixed to the result obtained from the 2D model.
In Figure 3, we present the 2D structure function, where the lines indicate the projections along the major and minor axes of the fitted model described in Equation 4. Our analysis shows that the degree of anisotropy, defined as the ratio of the major to minor axis of the diffractive scale, is approximately 2.17. In many cases, the anisotropy is aligned with the Earth’s magnetic field, similar to the large field-aligned structures reported by Loi et al. (2015) in MWA observations. When the anisotropy is not aligned with the magnetic field, it might instead be caused by travelling ionospheric waves (TIDs), where the smallest diffractive scales are observed along the direction of wave propagation (Mevius et al., 2016; Banerjee et al., 2025).
3.2 Anisotropy in the ionospheric structures
In this section, we studied how uneven (or anisotropic) the ionosphere is, and what direction the longest part of that unevenness (the major axis) points, compared to north. Previous studies (Spencer, 1955; Singleton, 1970; Wheelon, 2001; Loi et al., 2015; Mevius et al., 2016) found that irregularities in the ionosphere tend to stretch along Earth’s magnetic field lines. This alignment is particularly expected in the case of field-aligned irregularities or elongated plasma structures, which can be guided by the geomagnetic field. However, such alignment is not universal and may be disrupted by large-scale wave-like disturbances such as Travelling Ionospheric Disturbances (TIDs) (Mevius et al., 2016).
To investigate this potential anisotropy, we computed the Earth’s magnetic field vector at the GMRT site for the date of our observation (23rd November, 2024) using the World Magnetic Model (Chulliat et al., 2020). The magnetic field was evaluated at an altitude of approximately 300 km, which is a typical average height for peak ionospheric electron density according to the IRI model. Since the angle between our observing baselines and the magnetic field lines changes over time, we can’t just calculate one value of variation for each baseline. Instead, we grouped the data by both angle and length.
Figure 4: The figure shows the phase structure function where the data is binned by the angle relative to the projected Earth’s magnetic field. The colour bar indicates the angle in degrees.
Figure 4 displays the phase structure function, with the phase variance data grouped based on the angle relative to the projected Earth’s magnetic field. The angle between each baseline vector and the local magnetic field was determined using:
$\cos(\theta)=\frac{\mathbf{b}\cdot\mathbf{B}}{|\mathbf{b}||\mathbf{B}|},$
where $\mathbf{b}$ represent the baseline vector and $\mathbf{B}$ denotes the magnetic field vector.
We notice the phase variance varies for each a fixed length baseline when the angle between the baselines and projected magnetic field changes. For the lowest and highest angle bins this variation is close to two orders of magnitude in the phase variance (Figure 4). About $60\%$ of the LOFAR observations reported in Mevius et al. (2016) find the ionospheric structures are aligned with the magnetic field. This effect might be caused by ”density ducts”, long, tube-shaped features aligned with the field lines, as seen in Loi et al. (2015).
For radio interferometric calibration, the difference in signal behaviour along and across ionospheric structures needs to be considered when calculating noise levels. Additionally, ionospheric models used for calibration such as SPAM (Intema, H. T. et al., 2009) could be improved by incorporating the fact that these structures tend to follow the magnetic field and are tilted relative to the Earth’s surface (Mevius et al., 2016).
To explore this pattern more closely in our observation, we focused on two specific angle ranges: $0^{\circ}$$20^{\circ}$, which is approximately aligned with the magnetic field, and $70^{\circ}$$90^{\circ}$, which is nearly perpendicular to it. The 1D structure function model fits (Eqn. 2) show the diffractive scale along the field lines is much greater than the perpendicular direction (Figure 5). This suggest the ionospheric structures in our observation is not filed aligned. If we choose a subset of baselines above 1 km (Figure 6) the model fit improves (The Residual Standard Error (RSE) is reduced by $\sim 70\%$ and $\sim 43\%$ for the aligned and perpendicular fits, respectively) but our conclusion remain unchanged. We find the slope of the power law is close to 2. A possible cause of this anisotropy is medium-sized Travelling Ionospheric Disturbances (MSTIDs) (Vanvelthoven, 1990). These are wave-like patterns in the ionosphere that are usually not aligned with the magnetic field. They could explain cases where we see anisotropy, but not in the direction of the field.
Figure 5: Phase structure function comparing two groups of baselines based on their angle to the Earth’s magnetic field: those nearly aligned (between $0^{\circ}$ and $20^{\circ}$, shown with blue solid lines) and those nearly perpendicular (between $70^{\circ}$ and $90^{\circ}$, shown with red dotted lines). Each line shows a 1D power-law fit of phase variance with baseline length.
When we select a subset of baselines longer than 1 km (Figure 6), the model fit improves, but our conclusion remains unchanged. We find that the slope of the power law is close to 2. One possible cause of this anisotropy is medium-sized Travelling Ionospheric Disturbances (mTIDs) (Vanvelthoven, 1990). These are wave-like patterns in the ionosphere that are generally not aligned with the magnetic field and could explain the observed anisotropy that does not follow the field direction.
Figure 6: Same as Figure 5, but using only baselines longer than 1 km. The phase variance shows reduced scatter, resulting in a more reliable power-law fit.
4 Summary and Conclusions
In this study, we analysed nighttime ionospheric phase fluctuations using a 10-hour observation of the bright flux calibrator source 3C48 with the upgraded Giant Metrewave Radio Telescope (uGMRT). Specifically, we studied the ionospheric phase structure function at around 575-725 MHz while tracking the bright radio source 3C48. This function helps us understand how the phase of a radio signal changes with distance, due to turbulence in the ionosphere.
We measured differences in Total Electron Content (TEC) between pairs of antennas and converted them into phase delays. These measurements showed how the ionosphere varies across space and time. By analysing the phase differences as a function of distance between antennas, we created a structure function that shows clear power-law behaviour - a signature of turbulence. The best-fit slope of this function was slightly steeper than expected for classic Kolmogorov turbulence, which may indicate additional structures in the ionosphere.
We also found a diffractive scale of about 6.7 km, meaning the signal’s phase becomes significantly different across distances larger than this. A smaller diffractive scale suggests the ionosphere is less stable, causing faster and more complex changes in the signal. This has practical implications: to get clear radio images, we need to calibrate the data more frequently and in many directions, especially when the ionosphere is active.
While some other studies (e.g., using LOFAR) observed flattening at short distances - likely due to calibration noise - we did not see this in our data. For now, we kept our model simple, assuming a pure power-law without adding any extra noise component.
Further, we extended our analysis of the ionospheric phase structure function to two dimensions, allowing us to investigate directional effects. While the slope of the structure function ($\beta$) remains approximately constant across different directions, the diffractive scale $r_{\text{diff}}$ shows clear directional dependence. This indicates that the ionospheric turbulence is anisotropic.
We observed a characteristic band-like pattern in the 2D structure function, similar to previous LOFAR studies (e.g., Mevius et al. 2016). This pattern is probably due to large-scale, wave-like features in the ionosphere — such as Medium-Scale Travelling Ionospheric Disturbances (MSTIDs) - that move in the direction where the smallest diffractive scales are seen.
To model this anisotropy, we fitted the structure function using a 2D power-law model. The fit yielded an anisotropy ratio (major-to-minor axis of the diffractive scale) of approximately 2.17. In many cases, the direction of anisotropy aligns with the Earth’s magnetic field, consistent with previous observations of field-aligned structures (Loi et al., 2015). When misaligned, the anisotropy may instead reflect the propagation direction of ionospheric waves.
To further assess anisotropy in the phase fluctuations, we compared the phase structure function along baselines that are approximately aligned with the Earth’s magnetic field (0–20) and those nearly perpendicular to it (70–90). Our results show that the diffractive scale is significantly larger in the aligned direction than in the perpendicular one (Figure 5). This indicates that the ionospheric irregularities in our observation are not aligned with the magnetic field lines.
When we repeat the analysis using only long baselines (greater than 1 km), the fits become more stable, but the conclusion remains the same (Figure 6). The slope of the power-law structure function in both directions is close to 2 within $2\sigma$ error bars, which is steeper than the typical Kolmogorov value. These observations suggest the presence of medium-scale Travelling Ionospheric Disturbances (MSTIDs), which are known to produce wave-like structures that are not necessarily aligned with the magnetic field (Vanvelthoven, 1990). Therefore, the anisotropy we observe is likely caused by these dynamic ionospheric features, rather than static, field-aligned structures.
It is important to note that this study used only nighttime observations taken during the winter season. Therefore, our results do not account for diurnal or seasonal variations. Overall, our results demonstrate the capability of the uGMRT for ionospheric studies at mid-latitudes close to the geomagnetic equator (Opio et al., 2015).
Acknowledgements
We thank the staff of the GMRT for their support in making these observations possible. The GMRT is operated by the National Centre for Radio Astrophysics (NCRA) of the Tata Institute of Fundamental Research (TIFR). The authors utilized ChatGPT for AI-assisted copy editing and improving the manuscript’s language. AG would like to thank IUCAA, Pune, for providing support through the associateship programme, including access to the computational facility at IUCAA.
All authors acknowledge the financial support received through the SERB-SURE grant (SUR/2022/000595) from the Science and Engineering Research Board, a statutory body under the Department of Science and Technology (DST), Government of India.
Appendix A Sub-Band Variations in Ionospheric Phase Structure Function (Band-4)
In this appendix, we present additional structure function analyses at different sub-bands within the uGMRT Band-4 frequency range to assess the frequency dependence of ionospheric phase fluctuations. Figures 7, 8, and 9 illustrate the results for two representative frequencies: 612.5 MHz and 712.5 MHz. These include the one-dimensional phase structure function with power-law fits, two-dimensional anisotropic structure functions highlighting directional differences in turbulence, and binned phase variance plots as a function of baseline orientation relative to the projected Earth’s magnetic field. Together, these plots provide insights into how the ionospheric irregularities and anisotropic characteristics evolve across Band-4 frequencies.
Figure 7: Phase structure function as a function of baseline length. Left: At 612.5 MHz. Right: At 712.5 MHz. Blue points show the observed phase variance, and the red lines indicate the fitted power-law model.
Figure 8: Two-dimensional phase structure function showing anisotropy. Left: At 612.5 MHz. Right: At 712.5 MHz. The green dotted and red solid lines indicate the fitted major and minor diffractive axes, respectively, based on the anisotropic structure function model.
Figure 9: Phase structure function binned by angle relative to the projected Earth magnetic field. Left: At 612.5 MHz. Right: At 712.5 MHz. The colour bar indicates the angle in degrees.